# Motion Capture from Inertial and Vision Sensors

### v1

@@ Abstract
Human motion capture is the foundation for many computer vision and graphics tasks. While industrial motion captures systems with complex camera arrays or expensive wearable sensors have been widely adopted in movie and game production, consumer-affordable and easy-to-use solutions for personal applications are still far from mature. To utilize a mixture of a monocular camera and very few inertial measurement units (IMUs) for accurate multi-modal human motion capture in daily life, we contribute MINIONS in this paper, a large-scale Motion capture dataset collected from INertial and visION Sensors. MINIONS has several featured properties: 1) large scale of over five million frames and 400 minutes duration; 2) multi-modality data of IMUs signals and RGB videos labeled with joint positions, joint rotations, SMPL parameters, etc.; 3) a diverse set of 146 fine-grained single and interactive actions with textual descriptions. With the proposed MINIONS, we conduct experiments on multi-modal motion capture and explore the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs. The experiment results emphasize the unique advantages of inertial and vision sensors, showcasing the promise of consumer-affordable multi-modal motion capture and providing a valuable resource for further research and development.
Keywords:
Multi-modal Motion Capture Dataset and Benchmark
1 Introduction
@@
Table 1: Comparisons of publicly available human motion datasets for motion capture. MINIONS has more types of actions (Act), more frames, and a longer duration (minutes) of videos with single (S) and multiple (M) actors. Moreover, MINIONS provides both full-body IMU data and HD RGB videos annotated with 3D SMPL Mesh of actors, joints (KP), fine-grained action labels (F-Act), and texture of actors (Texture).
@@
MINIONS (Ours) 2024 146 5.5M 440 M
To this end, we contribute a large-scale Motion capture dataset from INertial and visION Sensors, named MINIONS, as shown in Figure 1. The MINIONS dataset has several featured properties: 1) Multi-modalities: it is collected from eight 2K cameras and suits with 17 nine-axis IMU sensors. An RGB-D scanner is also adopted to obtain a textured mesh for each actor. 2) Scalibility: it contains over 5.5 million frames and 440 minutes of action sequences captured from different viewpoints. 3) Diversity: it covers 146 categories of fine-grained actions performed by 36 groups of actors (20 actors for single actions and 16 groups of actors for multi-person interactions). 4) Abundance: it provides abundant time-synchronized annotations for each frame, including 2D/3D joints, SMPL parameters [31], fine-grained action class, and texture of actors. We compare our MINIONS with existing motion capture datasets in Table 1.
Based on MINIONS, we propose a baseline analysis on multi-modal human motion capture with both inertial and vision sensors, and validate the feasibility of stable motion capture using a monocular camera and very few IMUs. More specifically, our experiments reveal that monocular motion capture is often limited by jittering due to blurriness or occlusion, while IMUs-based motion capture tends to suffer from global position drifts caused by casual daily actions and loose-fitting clothes. However, a setup with four to six IMUs and a monocular camera can effectively achieve stable and virtually drift-free motion capture. Our baseline analysis offers substantial experimental evidence that paves the way for further research in this field. Furthermore, MINIONS can also be a benchmark dataset for many other tasks such as 2D-to-3D pose estimation [63, 40], fine-grained action recognition [15, 55], etc. The evaluation results of these tasks further exploit the potential of our dataset.
In summary, the contributions of this paper are three-fold:
  • We conduct a detailed baseline analysis of multi-modal human motion capture using both inertial and vision sensors, and further explore the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs.
2 Related Work
Motion Capture by Inertial Sensors [56, 22]. IMUs-based human motion capture systems have been widely applied in the industry because they are robust to changes in illumination and occlusion. Industrial solutions usually exploit full-body IMU sensors which must be bound to human bodies. For example, the Perception Neuron system [1] and the Xsens MVN system [42] both use 17 IMUs for human motion capture. Existing datasets for IMU-based motion capture, e.g., DIP-IMU [20], are collected with such systems. However, the expensive cost, e.g., thousands of dollars for the IMU suit, and complex setup procedures before motion capture greatly raise the application threshold. To reduce the cost of devices and setup, researchers study methods using very few IMUs such as six IMUs. Early studies focused on search-based methods [41, 47]. These methods first recorded the parameters of template actions to build a database. During the inference stage, they used a lazy learning strategy [5] to search for similar actions in the database. However, due to the weak correlation between human motions and sensor parameters, their accuracy and stability are far from applications. Recent approaches mainly adopted optimization-based or regression-based paradigms, which used machine learning or deep learning techniques. For example, optimization-based methods [35, 14] optimized all pose parameters in a sequence to find an optimal pose trajectory that is consistent with the acceleration parameters of sensors. Regression-based methods [20, 57], utilized deep neural networks to learn a mapping from sensor measurements to body positions and joint rotations from pair-wised data. These learning-based methods provide a promising solution for sparse IMUs-based human motion capture. However, they still suffer from location drift due to the cumulative error of IMUs over a long time, e.g., several minutes. Therefore, we consider visual information from RGB videos as supplementary cues for long-term stable motion capture.
Motion Capture from Monocular Video [25, 60, 61]. Motion capture, i.e., 3D pose estimation, from monocular videos has been a popular topic in the computer vision community [30]. Existing methods mainly utilize parameterized models like SCAPE [7], SMPL [31], and GHUM [6] as an intermediate representation of human motion. Motion capture from videos is formulated as the estimation of model parameters, i.e., positions of joints, rotations of joints, and shapes of bodies, from RGB frames. Recent approaches can also be divided into optimization-based methods and regression-based ones. Optimization-based methods [10, 17, 45] optimized parameters of the model, e.g., SMPL, to fit the joints or silhouettes of human bodies in video frames. These methods usually relied on the accuracy of the joint detector while inevitably suffering from the trade-off between prediction accuracy and computation cost. Regression-based methods, e.g., HMR [24] and ROMP [46], learned deep neural networks from large-scale data to regress the model parameters given a single RGB image. Popular datasets for training these methods include Human3.6M [21], MPI-INF-3DHP [37], 3DPW [33], etc., as listed in Table 1. Although these methods have achieved excellent accuracy and real-time performance on these benchmarks, they may still generate poor results and temporal jitters due to occlusion, fast actions, and subtle movements in real-world scenarios. Therefore, this paper aims to exploit sparse IMUs to overcome these challenges.
@@ 2 Related Work
Motion Capture by Combination Schemes. Motion capture from both videos and IMUs has also attracted the attention of researchers. Existing methods [18, 32, 34, 27, 44] effectively improve the motion capture accuracy by eliminating multiplayer ambiguity in videos and minimization of location drift of IMUs. For example, Gilbert et al. [16] fused multi-channel volumetric data from multi-view cameras and IMU signals to estimate 3D joints. They also built a dataset, TotalCapture [50], as listed in Table 1. However, TotalCapture only contained 50-minute videos of five types of actions performed by five subjects, limiting the scalability of the dataset and related methods. During the data collection of TotalCapture, actors are mandated to wear tight-fitting attire, resulting in unavoidable discrepancies between the attire worn in daily life and that of the experimental setting. Therefore, we focus on building a large-scale motion capture dataset of daily life actions containing both RGB videos and IMUs and providing a more practical motion capture paradigm using monocular videos and sparse inertial signals for daily applications.
Figure 2: Overview of hardware setup. (a) Full-body Perception Neuron Studio [1] with 17 IMUs in total. (b) Scenarios larger than 36 square meters, using 4 to 8 synchronized cameras. (c) A professional RGB-D scanner for texture recovery.
3 The MINIONS Dataset
@@
The construction of the MINIONS dataset consists of 3.1 Hardware Setup, 3.2 Calibration, 3.3 Textured Mesh Reconstruction, and 3.4 Human Motion Capture.
3.1 Hardware Setup
@@
We collect raw data in multiple scenes using four to eight synchronized cameras and full-body IMU suits with $17$ sensors, as shown in Figure 2 (a)(b).
@@
IMUs: Perception Neuron Studio [1] is applied to acquire full-body inertial data including acceleration, angular velocity, and magnetic orientation. Note that the synchronization between all cameras and IMUs is dependent on the calibrated timecode devices.
RGB-D Scanner: A professional PUNE Scanner based on Microsoft’s Kinect [3], as shown in Figure 2 (c), is used to obtain the static human surface for texture and SMPL shape parameters recovery. The scanner can provide a scan accuracy of less than 1mm, with realistic texture maps of $1280\times 1024$ resolution.
Figure 3: Overview of dataset construction. (a) Textured mesh reconstruction with an RGB-D scanner; (b) 3D joints triangulation and tracking from multi-view videos; (c) Human pose from full-body IMUs data; and (d) Motion recovery from inertial and visual results.
3.2 Calibration
@@
Calibration of IMUs. The raw sensor data are measured in the IMU local coordinate system and should be aligned with the SMPL coordinate system. So, we transform the raw data following the DIP approach [20]. Through T-pose calibration, as shown in Figure 2 (a), we obtain maps $R^{IS}$:$F^{S}\rightarrow{F^{I}}$, $R^{TI}$:$F^{I}\rightarrow{F^{T}}$ and $R^{TB}$:$F^{B}\rightarrow{F^{T}}$ between sensor coordinate system $F^{S}$, inertial coordinate system $F^{I}$, SMPL coordinate system $F^{T}$, and bones coordinate system $F^{B}$.
Calibration between Cameras and IMUs. By aligning the T-pose skeleton in the camera coordinate system $F^{C}$ and the SMPL coordinate system $F^{T}$, we obtain the transformation between them $R^{TC}$:$F^{C}\rightarrow{F^{T}}$.
3.3 Textured Mesh Reconstruction
As shown in Figure 3 (a), we obtain RGB-D images from the scanner and convert them to point clouds to generate human models $M_{t}$ with textures in a canonical pose [39, 58]. To obtain the static SMPL shape parameters $\beta$, we adjust the parameters $(\beta,\theta)$ of SMPL $M_{s}$ to be close to $M_{t}$ and in the interior of $M_{t}$ through the optimization as in [58],
@@ 3.3 Textured Mesh Reconstruction
where $\beta$ and $\theta$ denote the shape and pose parameters of SMPL, $E_{J}$ penalizes the error between the projection and the observed 2D joints, $E_{skin}$ keeps the scan points belonging to the skin close to the model $M_{s}$, $E_{cloth}$ prevents scan points belonging to clothes to be inside the model $M_{s}$, $E_{reg}$ is a priori term to make the results more reasonable.
3.4 Human Motion Capture
The human motion capture workflow includes 1) 2D joint detection; 2) 3D joint triangulation; 3) tracking; and 4) motion recovery. It provides high-quality annotations including 2D and 3D joints, SMPL parameters, and person identity (ID). To ensure the quality of the dataset, each step is double-checked by experienced annotators, and the erroneous results of detection and tracking are adjusted manually.
@@ 3.4 Human Motion Capture
3D Joints Triangulation. With the camera parameters $(K,R|T)$ and multi-view 2D joints $P_{2d}$, as shown in Figure 3 (b), we select top-$K$ views with the smallest reconstruction error to triangulate the initial 3D joints $\widetilde{P_{3d}}$, instead of all redundant camera views. To guarantee the quality of reconstruction, we discard the views with large reconstruction errors higher than $0.05$ meters.
Tracking. Unlike the workflow of single-subject motion datasets [21, 37, 50], our dataset track the 3D joints $\widetilde{P_{3d}}$ to get correct multi-person motion trajectories. To achieve efficient and accurate tracking, we estimate the corresponding 3D boxes from the 3D joints $\widetilde{P_{3d}}$ and then match these boxes using the SORT algorithm [8] with the max range as ten frames. Each actor has a distinguished ID, and the erroneous tracking results (including ID switches and assigning new IDs) are adjusted with manual double-checking.
Figure 5: Fine-grained Actions. MINIONS contains 121 single-player actions and 25 multi-player actions including common person-person and person-object interactive actions in daily life.
Motion Recovery. Most human motion datasets [21, 37] with single subject directly recover motion sequences from pure vision sensor. However, as shown in Figure 4 (a), occlusion between multiple subjects and objects is inevitable, a serious problem for pure visual recovery. Fortunately, despite the global position drift, we can still obtain accurate motions of actors wearing full-body IMUs under occlusion and varied illumination, as shown in Figure 3 (c) and Figure 4 (b). Therefore, for the joints that cannot be captured accurately by previous visual steps, we fill in the missing joints through full-body IMUs data to obtain the final 3D joints $P_{3d}$ and 3D mesh. In detail, for frames $x_{i}$ whose joints cannot be determined due to occlusions or low illumination, we search simultaneously forward and backward for visual frames $x_{i-n}$ and $x_{i+m}$ with high confidence ($n,m\geq 1$). Based on the high-confidence joints of $x_{i-n}$ and $x_{i+m}$, and IMUs data such as acceleration and velocity from $x_{i-n}$ to $x_{i+m}$, we estimate the ground truth joint locations of frame $x_{i}$. According to 3DPW [33], the accuracy of IMUs decreases as $m$ and $n$ increase over time. However, the values of $m$ and $n$ are typically controllable due to the aid of visual information, thus ensuring the accuracy of our estimation. After the fusion of inertial and visual information, we adopt SMPL [31] as the parameterized representation and fit it to the final 3D joints $P_{3d}$, as shown in Figure 3 (d) and Figure 4 (c). Similar to SMPLify [10], we constrain the SMPL parameters as follows:
$E(\beta,\theta,t)=E_{3d}+E_{reg},$ (2)
where $E_{3d}$ penalizes the error between the generated 3D joints and the final 3D joints $P_{3d}$, $E_{reg}$ is a priori term applied to make the results more reasonable. Specifically, the $E_{3d}$ loss is as follows:
@@
where $\beta$ denotes the static SMPL shape parameters obtained in 3.3, $\theta$, $t$, and $w_{3d}$ denote the SMPL pose parameters, translation of the camera, and confidence scores of detection, $\mathcal{J}$ is a pre-trained linear regression matrix used to generate 3D joints from the SMPL mesh $M(\theta,\beta,t)$.
@@
Figure 6: Qualitative results from single-subject motion capture data collection. Results of 3D mesh and the corresponding re-projected full-body 2D joints from various views.
3.5 Dataset Statistics
The detailed statistics of the MINIONS are listed in Table 1. Moreover, MINIONS is divided into MINIONS-S and MINIONS-M depending on the number of actors. MINIONS-S is collected for single-actor actions, which contain a total of 4.5 million video frames, 315 minutes of video duration, and 121 categories of fine-grained single actions. In contrast, MINIONS-M captures the interactive actions between multiple actors, containing a total of one million video frames, 125 minutes of video duration, and 25 categories of common person-person and person-object interactive actions in daily life.
The statistic on the number of frames for different action sequences reflects that most action sequences are smaller than 300 frames (6 seconds) and the longest one can be up to 1500 frames (50 seconds), which shows the complexity of human actions in life. We show some examples of our dataset in Figure 5. Please refer to the supplementary materials for more information on fine-grained actions.
@@ 3.5 Dataset Statistics
Figure 7: Qualitative results from multi-subjects motion capture data collection. Results of 3D mesh and the corresponding re-projected full-body 2D joints from various views. The identities of subjects are distinguished by colors.
4 Experiments
In this section, we first show the qualitative results of our dataset. Then we conduct a detailed baseline analysis of multi-modal human motion capture on our MINIONS and further explore the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs. At last, we evaluate the mainstream methods on various tasks, such as 2D-to-3D pose estimation and fine-grained video action recognition.
4.1 Qualitative Results

### v3

@@ Abstract
Human motion capture is the foundation for many computer vision and graphics tasks. While industrial motion capture systems with complex camera arrays or expensive wearable sensors have been widely adopted in movie and game production, consumer-affordable and easy-to-use solutions for personal applications are still far from mature. To utilize a mixture of a monocular camera and very few inertial measurement units (IMUs) for accurate multi-modal human motion capture in daily life, we contribute MINIONS in this paper, a large-scale Motion capture dataset collected from INertial and visION Sensors. MINIONS has several featured properties: 1) large scale of over five million frames and 400 minutes duration; 2) multi-modality data of IMUs signals and RGB videos labeled with joint positions, joint rotations, SMPL parameters, etc.; 3) a diverse set of 146 fine-grained single and interactive actions with textual descriptions. With the proposed MINIONS dataset, we propose a SparseNet framework to capture human motion from IMUs and videos by discovering their supplementary features and exploring the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs. The experiment results emphasize the unique advantages of inertial and vision sensors, showcasing the promise of consumer-affordable multi-modal motion capture and providing a valuable resource for further research and development.
I Introduction
@@
To this end, we contribute a large-scale Motion capture dataset from INertial and visION Sensors, named MINIONS, as shown in Figure 1. The MINIONS dataset has several featured properties: 1) Multi-modalities: it is collected from eight 2K cameras and suits with 17 nine-axis IMU sensors. An RGB-D scanner is also adopted to obtain a textured mesh for each actor. 2) Scalibility: it contains over 5.5 million frames and 440 minutes of action sequences captured from different viewpoints. 3) Diversity: it covers 146 categories of fine-grained actions performed by 36 groups of actors (20 actors for single actions and 16 groups of actors for multi-person interactions). 4) Abundance: it provides abundant time-synchronized annotations for each frame, including 2D/3D joints, SMPL parameters [38], fine-grained action class, and texture of actors. We compare our MINIONS with existing motion capture datasets in TABLE I.
Based on MINIONS, we propose a two-branch framework, named SparseNet, for motion capture by fusion of videos and IMUs. The core of SparseNet is a deliberately designed two-branch structure grounded in Bayesian probability theory and Matrix Fisher priors [13, 30]. For the sparse IMU signals, the Sparse IMUs branch takes inertial measurements as input to learn posture features, providing rotational probability distribution. To overcome the incompleteness and location drift of sparse IMUs, we design a visual branch to supplement the missed joints and correct the location drift, estimating 3D bone probability distribution and pose probability prior. This framework enables adaptive integration through a posterior fusion module that aligns distributions from both branches. Experiments reveal that the visual branch for monocular motion capture is often limited by jittering due to blurriness or occlusion, while the IMUs branch tends to suffer from global position drifts. However, a setup with four to six IMUs and a monocular camera can effectively achieve stable and virtually drift-free motion capture. Our experiment analysis offers substantial experimental evidence that paves the way for further research in this field. Furthermore, MINIONS can also be a benchmark dataset for many other tasks such as 2D-to-3D pose estimation [74, 47], fine-grained action recognition [18, 63], etc. The evaluation results of these tasks further exploit the potential of our dataset.
In summary, the contributions of this paper are three-fold:
  • We conduct SparseNet, a two-branch framework of multi-modal human motion capture using both inertial and vision sensors, and further explore the possibilities of consumer-affordable motion capture using a monocular camera and very few IMUs.
II Related Work
Motion Capture by Inertial Sensors [64, 27]. IMUs-based human motion capture systems have been widely applied in the industry because they are robust to changes in illumination and occlusion. Industrial solutions usually exploit full-body IMU sensors which must be bound to human bodies. For example, the Perception Neuron system [1] and the Xsens MVN system [49] both use 17 IMUs for human motion capture. However, the expensive cost, e.g., thousands of dollars for the IMU suit, and complex setup procedures before motion capture greatly raise the application threshold. To reduce the cost of devices and setup, researchers study methods using very few IMUs such as six IMUs. Recent approaches mainly adopted optimization-based or regression-based paradigms, which used machine learning or deep learning techniques. For example, optimization-based methods [41, 9] optimized all pose parameters in a sequence to find an optimal pose trajectory that is consistent with the acceleration parameters of sensors. Regression-based methods [25, 66], utilized deep neural networks to learn a mapping from sensor measurements to body positions and joint rotations from pair-wised data. These learning-based methods provide a promising solution for sparse IMUs-based human motion capture. However, they still suffer from location drift due to the cumulative error of IMUs over a long time, e.g., several minutes. Therefore, we consider visual information from RGB videos as supplementary cues for long-term stable motion capture.
Motion Capture from Monocular Video [31, 70, 71]. Motion capture, i.e., 3D pose estimation, from monocular videos has been a popular topic in the computer vision community [37]. Recent approaches [19, 11] can be divided into optimization-based methods and regression-based ones. Optimization-based methods [5, 21, 52] optimized parameters of the model, e.g., SMPL, to fit the joints or silhouettes of human bodies in video frames. Regression-based methods, e.g., HMR [29] and ROMP [53], learned deep neural networks from large-scale data to regress the model parameters given a single RGB image. Although these methods have achieved excellent accuracy and real-time performance on these benchmarks, they may still generate poor results and temporal jitters due to occlusion, fast actions, and subtle movements in real-world scenarios. Therefore, this paper aims to exploit sparse IMUs as supplementary cues to overcome these challenges.
@@ II Related Work
TABLE I: Comparisons of publicly available human motion datasets for motion capture. MINIONS has more types of actions (Act), more frames, and a longer duration (minutes) of videos with single (S) and multiple (M) actors. Moreover, MINIONS provides both full-body IMU data and HD third-person view (TPV) RGB videos annotated with 3D SMPL Mesh of actors, joints (KP), fine-grained action labels (F-Act), and texture of actors (Texture).
@@
RELI11D [62] 2024 15 0.2M 199 M
EMHI [16] 2025 39 3.1M - S
MINIONS (Ours) - 146 5.5M 440 M
@@
Motion Capture by Combination Schemes [10]. Motion capture from both videos and IMUs has also attracted the attention of researchers. Existing methods [23, 33, 51] effectively improve the motion capture accuracy by eliminating multiplayer ambiguity in videos and minimization of location drift of IMUs. For example, Gilbert et al. [20] fused multi-channel volumetric data from multi-view cameras and IMU signals to estimate 3D joints. They also built a dataset, TotalCapture [56], as listed in TABLE I. However, TotalCapture only contained 50-minute videos of five types of actions performed by five subjects, limiting the scalability of the dataset and related methods. During the data collection of TotalCapture, actors are mandated to wear tight-fitting attire, resulting in unavoidable discrepancies between the attire worn in daily life and that of the experimental setting. Therefore, we focus on building a large-scale motion capture dataset of daily life actions containing both RGB videos and IMUs and providing a more practical motion capture paradigm using monocular videos and sparse inertial signals for daily applications.
III The MINIONS Dataset
@@
The construction of our proposed MINIONS dataset consists of III-A Multimedia Hardware Setup, III-B Calibration, III-C Textured Mesh Reconstruction,  III-D Global Motion Annotations, and  III-E Dataset Statistics.
III-A Multimedia Hardware Setup
@@
As shown in Fig. 2 (a), we collect raw data in multiple scenes with multimedia hardware, including a professional RGB-D scanner, four to eight synchronized cameras, and full-body IMU suits with $17$ sensors.
RGB-D Scanner: A professional PUNE Scanner based on Microsoft’s Kinect [3], is used to obtain the human surface for texture and SMPL shape parameters recovery. The scanner can provide a scan accuracy of less than 1mm, with realistic texture maps of $1280\times 1024$ resolution.
@@
IMUs: The Perception Neuron Studio [1] is applied to acquire full-body inertial data, including acceleration, angular velocity, and magnetic orientation. Note that the synchronization between all cameras and IMUs is dependent on the calibrated timecode devices.
Figure 2: Overview of Dataset Construction. (a) Input multimedia data collected from multimedia hardware, including the RGB-D scanner, multiple synchronized cameras, and full-body IMU suits; (b) Data pre-processing, including point cloud data conversion, 2D joint detection, 3D joints triangulation, and IMUs data alignment; (c) Textured 3D human models and approximated SMPL shape; and (d) Global motion recovery from inertial and visual results.
III-B Calibration
@@
Calibration of IMUs. The raw sensor data are measured in the IMU local coordinate system and should be aligned with the SMPL coordinate system. So, we transform the raw data following the DIP approach [25]. Through initial T-pose calibration, we obtain maps $R^{IS}$$F^{S}\rightarrow{F^{I}}$, $R^{TI}$$F^{I}\rightarrow{F^{T}}$ and $R^{TB}$$F^{B}\rightarrow{F^{T}}$ between sensor coordinate system $F^{S}$, inertial coordinate system $F^{I}$, SMPL coordinate system $F^{T}$, and bones coordinate system $F^{B}$. The sensor coordinate system $F^{S}$ is the local coordinate system attached to each IMU sensor, moving with the sensor. The inertial coordinate system $F^{I}$ is a fixed global reference coordinate system, often aligned with gravity and magnetic north at the start of calibration. The SMPL Coordinate System $F^{T}$ is the canonical coordinate system of the SMPL human body model, with its origin typically at the pelvis. The Bones Coordinate System $F^{B}$ is the local coordinate system attached to an individual bone or joint, defining its local rotation.
Calibration between Cameras and IMUs. By aligning the T-pose skeleton in the camera coordinate system $F^{C}$ and the SMPL coordinate system $F^{T}$, we obtain the transformation between them $R^{TC}$$F^{C}\rightarrow{F^{T}}$.
III-C Textured Mesh Reconstruction
As shown in the upper part of Figure 2 (b) and Figure 2 (c), we obtain RGB-D images from the scanner and convert them to point clouds to generate human models $M_{t}$ with textures in a canonical pose [46, 68]. To obtain the static SMPL shape parameters $\beta$, we adjust the parameters $(\beta,\theta)$ of SMPL $M_{s}$ to be close to $M_{t}$ and in the interior of $M_{t}$ through the optimization as in [68],
@@ III-C Textured Mesh Reconstruction
where $\beta$ and $\theta$ denote the shape and canonical pose parameters of SMPL, $E_{J}$ penalizes the error between the projection and the observed 2D joints, $E_{skin}$ keeps the scan points belonging to the skin close to the model $M_{s}$, $E_{cloth}$ prevents scan points belonging to clothes to be inside the model $M_{s}$, $E_{reg}$ is a priori term to make the results more reasonable.
III-D Global Motion Annotations
The global human motion annotations workflow includes 1) tracking; 2) 2D joint detection; 3) 3D joints triangulation, and 4) motion recovery from multimedia data. It provides high-quality annotations, including 2D and 3D joints, SMPL parameters, and person identity (ID). To ensure the quality of the dataset, each step is double-checked by experienced annotators, and the erroneous results of detection and tracking are adjusted manually.
Tracking. Unlike the workflow of single-subject motion datasets [26, 43, 56], our dataset track the 3D joints $\widetilde{P_{3d}}$ to get correct multi-person motion trajectories. To achieve efficient and accurate tracking, we estimate the corresponding 3D boxes from the 3D joints $\widetilde{P_{3d}}$ and then match these boxes using the reid-based [22] StrongSORT algorithm [14]. Each actor has a distinguished ID, and the erroneous tracking results (including ID switches and assigning new IDs) are adjusted with manual double-checking.
3D Joints Triangulation. With the camera parameters $(K,R|T)$ and multi-view 2D joints $P_{2d}$, as shown in Figure 2 (b), we select top-$K$ views with the smallest reconstruction error to triangulate the initial 3D joints $\widetilde{P_{3d}}$, instead of all redundant camera views. To guarantee the quality of reconstruction, we discard the views with large reconstruction errors higher than $0.01$ meters.
@@ III-D Global Motion Annotations
Figure 4: Fine-grained Actions. MINIONS contains 121 single-player actions and 25 multi-player actions, including common person-person and person-object interactive actions in daily life.
Motion Recovery from Multimedia data. Most human motion datasets [26, 43] with a single subject directly recover motion sequences from pure vision sensors. However, as shown in Figure 3 (a), occlusion between multiple subjects and objects is inevitable, a serious problem for pure visual recovery. Fortunately, despite the global position drift, we can still obtain accurate motions of actors wearing full-body IMUs under occlusion and varied illumination, as shown in Figure 3 (b). Therefore, for the joints that cannot be captured accurately by previous visual steps, we fill in the missing joints through full-body IMUs data to obtain the final 3D joints $P_{3d}$ and 3D mesh. In detail, for frames $x_{i}$ whose joints cannot be determined due to occlusions or low illumination, we search simultaneously forward and backward for visual frames $x_{i-n}$ and $x_{i+m}$ with high confidence ($n,m\geq 1$). Based on the high-confidence joints of $x_{i-n}$ and $x_{i+m}$, and IMUs data such as acceleration and velocity from $x_{i-n}$ to $x_{i+m}$, we estimate the ground truth joint locations of frame $x_{i}$. According to 3DPW [39], the accuracy of IMUs decreases as $m$ and $n$ increase over time. However, the values of $m$ and $n$ are typically controllable due to the aid of visual information, thus ensuring the accuracy of our estimation. After that, we adopt SMPL [38] as the parameterized representation and fit it to the 2D joints $P_{2d}$, 3D joints $P_{3d}$ and IMUs acceleration, as shown in Figure 2 (d). We constrain the global motion parameters as follows:
$E(\beta,\theta,t)=E_{2d}+E_{3d}+E_{smooth}+E_{reg},$ (2)
where $E_{2d}$ and $E_{3d}$ penalizes the error between the detected 2D/3D joints and the final 2D/3D joints, $E_{smooth}$ uses the acceleration information from the IMUs to smooth the final global motion, $E_{reg}$ is a priori term [5] applied to make the results more reasonable. The $E_{2d}$ and $E_{3d}$ loss is as follows:
$E_{2d}=\sum w_{2d}\cdot\left\|K\mathcal{J}M(\theta,\beta,t)-P_{2d}\right\|_{2}^{2},$ (3)
@@
where $\beta$ denotes the static SMPL shape parameters obtained in III-C, $\theta$, $t$, and $w_{2d}/w_{3d}$ denote the SMPL pose parameters, translation of the camera, and weighted scores, $\mathcal{J}$ is a pre-trained linear regression matrix used to generate 3D joints from the SMPL mesh $M(\theta,\beta,t)$, $K$ is the camera intrinsic parameters to transfer 3D joints to 2D joints.
@@
Figure 5: Qualitative results from single-subject motion capture collection. Results of 3D mesh and the corresponding re-projected full-body 2D joints from various views, visualized with white rendering for enhanced clarity.
@@
Figure 6: Qualitative results from multi-subject motion capture collection. Results of 3D mesh and the corresponding re-projected full-body 2D joints from various views. The identities of subjects are distinguished by colors.
III-E Dataset Statistics
The detailed statistics of the MINIONS are listed in TABLE I. MINIONS is divided into MINIONS-S and MINIONS-M depending on the number of actors. MINIONS-S is collected for single-actor actions, which contain a total of 4.5 million video frames, 315 minutes of video duration, and 121 categories of fine-grained single actions. In contrast, MINIONS-M captures the interactive actions between multiple actors, containing a total of one million video frames, 125 minutes of video duration, and 25 categories of common person-person and person-object interactive actions in daily life. The detailed action lists are shown in the supplementary materials. Note that informed consent was obtained from all individual participants included in the study.